# A statistical mechanical view of complex economies under the green transition

### v1

@@ V Nature of the phase transitions
In this Section, we provide some numerical evidence on the nature of the phase transitions illustrated in Fig. 1. Their origins can be traced back to Eq. (3), which effectively identifies a set of constraints. Namely, the economy’s production network matrix $q_{n}^{c}$ and the optimal scales of production at equilibrium $s_{n}^{\star}$ must guarantee that the production level of each good is a non-negative quantity, i.e., $(x^{c})^{\star}\geq 0$. Clearly, at finite $N$ and $C$, this depends on the specific draw of $q_{n}^{c}$ from the probability density in Eq. (10). It may very well be that, for some ‘unlucky’ draw of the input-output matrix, no production scale values exist to guarantee strictly positive consumption levels (at least for some goods), making the ‘shut down’ solution $s_{n}^{\star}=0$, $\forall\ n$ (and therefore $(x^{c})^{\star}=0$, $\forall\ n$) the only viable one.
In the thermodynamic limit $N,C\rightarrow\infty$, with $\nu=N/C$ fixed, the above process results in a sharp separation between regions where all draws of the input-output matrix are ‘unlucky’, resulting in systematic economic shut down, from regions where they all are ‘lucky’, resulting in some non-trivial level of economic activity. More formally, the former correspond to regions where the volume
$V=\int_{0}^{\infty}\mathrm{d}\bm{s}\prod_{c=1}^{C}\Theta\left(x_{0}^{c}+\sum_{n=1}^{N}s_{n}q_{n}^{c}\right)$ (14)
shrinks to zero in the thermodynamic limit.
We numerically corroborate the above picture by solving the following linear programming problem at finite $N$ and $C$ values
@@
We further corroborate the above picture by solving the linear programming problem in Eq. (15) — at fixed $x_{0}^{c}$ and $q_{n}^{c}$, i.e., for the same economy — for several different random choices of the vector $\rho$. For each economy, we build the correlation matrix of the solutions $\bm{s}^{\star}$ obtained in such a way, and in Fig. 5 we report the behavior of their leading eigenvalue $\ell_{\mathrm{MAX}}$, averaged over several different economies. As known from principal component analysis Livan et al. (2018), values $\ell_{\mathrm{MAX}}/N\rightarrow 1$ indicate very strong correlations, which in our case indicate that the solutions $\bm{s}^{\star}$ obtained for different random draws of $\bm{\rho}$ are exceedingly similar to each other, which, in turn, indicates that the volume in Eq. (15) must be shrinking. As it can be seen from Fig. 5, $\ell_{\mathrm{MAX}}/N$ displays a rather clear U-shaped behavior as a function of $\nu$, getting closer to one as $\nu$ approaches its two critical values ($\nu\approx 0.92$ and $\nu\approx 4.43$, respectively) for the parameters used in the simulation ($f_{0}=0.5$ and $\lambda=0.02$).
Figure 5: Average leading eigenvalue $\ell_{\mathrm{MAX}}$ (normalized by $N$) of the correlation matrix of different solutions $\bm{s}^{\star}$ of the linear programming problem in Eq. (15) obtained for different random choices of $\bm{\rho}$. We choose $N=150$, and fix $C=N/\nu$ accordingly (rounding up to the nearest integer). For each value of $\nu$, we generate 500 different economies (i.e., different draws of $x_{0}^{c}$ and $q_{n}^{c}$), using $f_{0}=0.5$ and $\lambda=0.02$. For each of those economies, we generate 100 different solutions of the linear programming problem, and use them to form the aforementioned correlation matrix.
Put together, the above numerical evidence is very much reminiscent of phase transitions in high-dimensional random geometries such as those found, e.g., in models of neural networks Gardner (1988) and financial markets Bardoscia et al. (2012), which have been characterized systematically in Donoho and Tanner (2009).
@@ VI Discussion
In this paper, we propose a stylized model of a complex economy to characterize the average behavior of a standard General Equilibrium Theory (GET) model when its parameters are treated as random variables, following the approach of the Statistical Mechanics of disordered systems. Our main focus is on the waste generated by complex economies, which we use as a proxy for emissions and other unwanted byproducts released by industrial production.
Our model is, of course, a simplified representation of the real-world complexity that characterizes modern economies and supply chains. To begin with, it inherits from GET the limitation of not capturing temporal dynamics, as GET focuses on static equilibria. Nonetheless, these equilibria can be interpreted in a ‘meta-stable’ sense — that is, as economic states that may persist for extended periods before transitioning to a new equilibrium due to shifts in the economy’s fundamental parameters.
Within this framework, our model serves as a useful tool for evaluating the tradeoffs imposed on modern economies by the so called green transition. In this context, the model’s phase diagrams (Fig. 1) illustrate the possible trajectories an economy may follow during the course of this transition. The central trade-off highlighted by our model lies between economic development and economic welfare, very much in line with the literature on ‘green growth’ and its alternatives D’Alessandro et al. (2020); Hickel and Kallis (2020). Specifically, we consistently find that more wasteful (i.e., brown) economies have access to higher values of $\nu=N/C$ — that is, a greater number of available technologies per good produced. However, much of this technological potential remains unrealized, as these economies exhibit a monotonic decline in the fraction $\phi$ of active technologies with increasing $\nu$. At the same time, greener economies, despite having access to lower values of $\nu$, tend to generate higher economic welfare, reflected in higher production scales, higher GDP, and higher consumption levels (see Fig. 2).
One of the most notable features of our model is that it reveals a double phase transition, separating the region where an economy can function from two regions of economic shutdown. These shutdown regions correspond to scenarios where either economic development is too advanced to be compatible with low levels of waste, or too limited for the economy to function. Interestingly, greener economies tend to approach the former boundary along a path marked by substantial growth, particularly in terms of operation scale (see, e.g., the top-left panel in Fig. 1). This suggests that rapid growth under low-waste conditions may require careful management to prevent abrupt disruptions in economic activity.
On a more technical note, we were able to explain the aforementioned phase transitions with rather convincing numerical evidence, linking them with the ‘phase space shrinkage’ that occurs as an economy is subjected to an increasing number of constraints. This behavior is reminiscent of algorithmic phase transitions commonly studied in Computer Science problems Monasson et al. (1999); Mézard et al. (2002).
Appendix A The optimization problem
@@
×$\left.\times\exp\left[\beta\sum_{a=1}^{r}U_{N}\left(\bm{x}_{a}\right)+\mathrm{i}\sum_{a=1}^{r}\sum_{c=1}^{C}\hat{x}_{a}^{c}\left(x_{a}^{c}-x_{0}^{c}\right)\right]\right.$
×$\left.\times\exp\left[-\frac{1}{2C}\sum_{n=1}^{N}\sum_{c=1}^{C}\left(\hat{z}_{n}-\sum_{a=1}^{r}s_{n,a}\hat{x}_{a}^{c}\right)^{2}+\mathrm{i}\sqrt{\lambda}w\sum_{n=1}^{N}\hat{z_{n}}+\mathrm{i}\epsilon_{0}\sum_{n=1}^{N}\hat{z}_{n}\right]\right\rangle_{w}\ ,$
where we have made use of the Hubbard-Stratonovich integral and introduced a Gaussian standard variable $w$, i.e., $w\sim\mathcal{N}(0,1)$. The factor in front of the average with respect to $w$ comes from the denominator in Eq. (10).
@@
11 11 $=$ $N\int_{-\infty}^{\infty}\mathrm{d}\omega_{ab}\ \delta\left(N\omega_{ab}-\sum_{n=1}^{N}s_{n,a}s_{n,b}\right)=N\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega_{ab}\mathrm{d}\hat{\omega}_{ab}}{2\pi\mathrm{i}}\exp\left[\hat{\omega}_{ab}\left(N\omega_{ab}-\sum_{n=1}^{N}s_{n,a}s_{n,b}\right)\right]$
11 11 $=$ $N\int_{-\infty}^{\infty}\mathrm{d}k_{a}\ \delta\left(Nk_{a}-\sum_{n=1}^{N}\hat{z}_{n}s_{n,a}\right)=N\int_{-\infty}^{\infty}\frac{\mathrm{d}k_{a}\mathrm{d}\hat{k}_{a}}{2\pi\mathrm{i}}\exp\left[\hat{k}_{a}\left(Nk_{a}-\sum_{n=1}^{N}\hat{z}_{n}s_{n,a}\right)\right]\ ,$
@@
$h\left(\Omega,\kappa,\hat{\kappa},\hat{\gamma},\chi,\hat{\chi}\right)$ $=$ $\frac{1}{2}\left(\Omega\hat{\chi}-\frac{\hat{\gamma}\chi}{\nu}\right)-\frac{\hat{\kappa}\kappa}{\nu}+\left\langle\max_{s\geq 0}\left[-\frac{\hat{\chi}}{2}s^{2}+st\sqrt{\hat{\gamma}-\left(1+\lambda\right)\hat{\kappa}^{2}}+\hat{\kappa}\epsilon_{0}s\right]\right\rangle_{t}$
@@
$h\left(\Omega,\kappa,\hat{\kappa},\hat{\gamma},\chi,\hat{\chi}\right)$ $=$ $\frac{1}{2}\left(\Omega\hat{\chi}-\frac{\hat{\gamma}\chi}{\nu}\right)-\frac{\hat{\kappa}\kappa}{\nu}-\frac{\hat{\chi}}{2}\left\langle(s^{\star})^{2}\right\rangle_{t}+\left\langle s^{\star}t\right\rangle_{t}\sqrt{\hat{\gamma}-\left(1+\lambda\right)\hat{\kappa}^{2}}+\hat{\kappa}\epsilon_{0}\left\langle s^{\star}\right\rangle_{t}$
@@
At last, posing $p=-\hat{\kappa}$ and $\sigma=\sqrt{\hat{\gamma}-\left(1+\lambda\right)\hat{\kappa}^{2}}$, the expression in Eq. (A) matches the one in Eq. (III). The saddle point condition (see Eq. (13)) finally provides us with a set of equations for the order parameters that fully specify the solution of our optimization problem:
$\Omega$ $=$ $\left\langle\left(s^{\star}\right)^{2}\right\rangle_{t}$
@@
Appendix B Statistical properties at equilibrium
In this section we derive the statistical properties of the optimal scales of production $s^{\star}$ and consumption $x^{\star}$ at equilibrium. Starting from the former, we see from the third term in Eq. (A) that it must be the solution of the following equation

### v2

@@ V Nature of the phase transitions
In this Section, we provide some numerical evidence on the nature of the phase transitions illustrated in Fig. 1. Their origins can be traced back to Eq. (3), which effectively identifies a set of constraints. Namely, the economy’s production network matrix $q_{n}^{c}$ and the optimal scales of production at equilibrium $s_{n}^{\star}$ must guarantee that the production level of each good is a non-negative quantity, i.e., $(x^{c})^{\star}\geq 0$. Clearly, at finite $N$ and $C$, this depends on the specific draw of $q_{n}^{c}$ from the probability density in Eq. (8). It may very well be that, for some ‘unlucky’ draw of the input-output matrix, no production scale values exist to guarantee non-negative consumption levels for all goods, making the ‘shut down’ solution $s_{n}^{\star}=0$, $\forall\ n$ (and therefore $(x^{c})^{\star}=0$, $\forall\ n$) the only viable one.
Remarkably, some analytical considerations can be made in order to characterize the above statistical process from a more formal point of view. To that end, let us introduce the following quantity:
$V=\int_{0}^{\infty}\mathrm{d}\bm{s}\prod_{c=1}^{C}\Theta\left(x_{0}^{c}+\sum_{n=1}^{N}s_{n}q_{n}^{c}\right)\ .$ (10)
In the above expression, we are integrating over all possible production scales—for fixed initial endowments and production matrix—a simple indicator function constructed as a product of Heaviside functions, i.e., $\Theta(x)=1$ for $x\geq 0$ and $\Theta(x)=0$ otherwise. Should our draw of the production matrix $q_{n}^{c}$ be particularly ‘unlucky’, it will be impossible to find even a single vector $\bm{s}$ that prevents at least one of the $C$ Heaviside functions from becoming equal to zero, resulting in $V=0$. Of course, we are not just interested in single instances of the production matrix, but rather on the average behavior of the quantity in Eq. (10), which effectively quantifies the ‘volume’ in the space of production scales where economic activity can take place, in the thermodynamic limit $N,C\rightarrow\infty$, with $\nu=N/C$ fixed. Specifically, we are interested in determining under what conditions, and for which parameter values, we can expect the above volume to shrink to zero.
We numerically address the above problem by solving the following linear programming problem at finite $N$ and $C$ values
@@
We further corroborate the above picture by solving the linear programming problem in Eq. (11) — at fixed $x_{0}^{c}$ and $q_{n}^{c}$, i.e., for the same economy — for several different random choices of the vector $\rho$. For each economy, we build the correlation matrix of the solutions $\bm{s}^{\star}$ obtained in such a way, and in Fig. 5 we report the behavior of their leading eigenvalue $\ell_{\mathrm{MAX}}$, averaged over several different economies. As known from principal component analysis [30], values $\ell_{\mathrm{MAX}}/N\rightarrow 1$ indicate very strong correlations, which in our case indicate that the solutions $\bm{s}^{\star}$ obtained for different random draws of $\bm{\rho}$ are exceedingly similar to each other, which, in turn, indicates that the volume in Eq. (11) must be shrinking. As it can be seen from Fig. 5, $\ell_{\mathrm{MAX}}/N$ displays a rather clear U-shaped behavior as a function of $\nu$, getting closer to one as $\nu$ approaches its two critical values ($\nu\approx 0.92$ and $\nu\approx 4.43$, respectively) for the parameters used in the simulation ($f_{0}=0.5$ and $\lambda=0.02$).
Put together, the above numerical evidence is very much reminiscent of phase transitions in high-dimensional random geometries such as those found, e.g., in models of neural networks [31] and financial markets [23], which have been characterized systematically in [32].
Figure 5: Average leading eigenvalue $\ell_{\mathrm{MAX}}$ (normalized by NN) of the correlation matrix of different solutions $\bm{s}^{\star}$ of the linear programming problem in Eq. (11) obtained for different random choices of $\bm{\rho}$. We choose $N=150$, and fix $C=N/\nu$ accordingly (rounding up to the nearest integer). For each value of $\nu$, we generate 500 different economies (i.e., different draws of $x_{0}^{c}$ and $q_{n}^{c}$), using $f_{0}=0.5$ and $\lambda=0.02$. For each of those economies, we generate 100 different solutions of the linear programming problem, and use them to form the aforementioned correlation matrix.
@@ VI Discussion
In this work, we developed a stylized model of a complex economy to analyze the trade-offs associated with the green transition, using techniques from the Statistical Mechanics of disordered systems. By promoting the parameters of a General Equilibrium Theory (GET) model to random variables, we characterized the typical, self-averaging properties of large random economies under environmental constraints, akin to those that modern developed economies are increasingly expected to face under the so called green transition. A central feature of our model is the explicit inclusion of a waste variable, which serves as a proxy for emissions or other undesirable byproducts of economic activity.
Our model necessarily makes a number of simplifying assumptions, and it inherits from GET the limitation of not capturing temporal dynamics, as GET focuses on static equilibria. Nonetheless, these equilibria can be interpreted in a ‘meta-stable’ sense—that is, as economic states that may persist for extended periods before transitioning to a new equilibrium due to shifts in the economy’s fundamental parameters. Despite its limitations, our model captures key qualitative features of modern economies under environmental pressure. The most notable outcome is the emergence of a double phase transition in the $(\nu,\lambda)$ plane, where $\nu=N/C$ denotes the ratio of technologies to goods (which proxies economic development) and $\lambda$ characterizes waste intensity. Economic activity is confined to a bounded ‘feasible’ region, surrounded by two shutdown regimes.
The first, at low $\nu$, corresponds to a technology-scarcity regime: the economy cannot sustain positive production because there are too few viable technological combinations to meet the basic non-negativity constraints on consumption (Eq. (3)). This regime is reminiscent of underdevelopment traps [33], where lack of industrial diversity prevents economic takeoff. The second, at high $\nu$, is a waste-constrained regime: while the economy possesses many technologies, the cumulative waste generated by activating them shrinks the feasible production space until it collapses. This reflects the tension between technological abundance and environmental constraints, and parallels the notion of planetary boundaries in sustainability studies [34].
Our results also highlight systematic differences between ‘brown’ and ‘green’ economies. Brown economies (high $\lambda$) can, in principle, access higher $\nu$ values, corresponding to a larger technological base, but they activate only a small fraction of their technologies as development progresses. This leads to lower production scales, GDP, and average consumption. Green economies (low $\lambda$), by contrast, operate with fewer technologies per good but utilize a higher fraction of them, achieving greater economic welfare across most of their feasible region. However, as green economies approach the high-$\nu$ boundary, they become vulnerable to abrupt shutdowns if environmental constraints tighten or if production scales are pushed too aggressively.
These findings offer an interpretation in standard economic terms. In classical GET, the production possibility frontier (PPF)—the boundary between what can and cannot be produced given a set of resources and technologies [35]—is smooth and convex. In our statistical mechanical formulation, the effective PPF can fragment and ultimately vanish as complexity and environmental pressure interact. The ‘volume collapse’ of feasible allocations that drives the double phase transition is analogous to the shrinking of solution spaces in high-dimensional combinatorial problems, a phenomenon well documented in the study of algorithmic phase transitions [36, 37]. From an economic perspective, this reflects the increasing difficulty of coordinating production in highly complex, environmentally constrained economies, and resonates with concerns about structural fragility in global value chains [2, 1].
Overall, our analysis provides a stylized yet insightful framework for thinking about the structural risks of the green transition. It shows that low-waste technological diffusion can initially enhance economic welfare by enabling higher activation of productive technologies. However, the feasible space for economic activity remains bounded, and rapid growth in a tightening environmental regime may lead to sudden collapses in output. These insights emphasize the importance of managing both technological development and environmental capacity in a coordinated manner to ensure a smooth transition to sustainable production paradigms.
Appendix A Statistical mechanical formulation
We are interested in solving the optimization problem in Eq. (5) as the economy grows large, i.e., when $N,C\rightarrow\infty$, with $\nu=N/C$ fixed. Thus, we are effectively interested in calculating the maxima of the utility function $U_{N}$ in the following limit:
where we have introduced the following function.
With this position, the optimization problem in Eq. (12) is effectively converted into a steepest descent search, as the limit $\beta\rightarrow\infty$ selects the values of the utility function $U_{N}$ that dominate the integral defining $Z_{N}$, which in turn provides us with the desired result. In Statistical Mechanics, $Z_{N}$ obviously mirrors a canonical partition function (and its logarithm mirrors a free energy), with the parameter $\beta$ playing the role of an inverse temperature. According to this interpretation, Eq. (12) informs us that the optimization problem we are interested in is fully equivalent to computing the zero temperature limit of the free energy of the system described by the partition function $Z_{N}$.
Reasonably, one can expect the maxima of the function $U_{N}$ to grow with the system’s size $N$, so that the quantity $U_{N}/N$ in Eq. (12) will remain finite (intensive, in statistical mechanical parlance) in the $N\rightarrow\infty$ limit. Thus, our position is that, for a reasonable choice of the utility function, the maxima of $U_{N}/N$ will cease to depend on the economy’s details as it grows large, i.e., they will obey
$\lim_{N\rightarrow\infty}\frac{1}{N}\max_{\bm{s}\geq 0}\ U_{N}(\bm{s}|\bm{q},\bm{\epsilon},\bm{x}_{0})=\lim_{N\rightarrow\infty}\frac{1}{N}\left\langle\max_{\bm{s}\geq 0}\ U_{N}(\bm{s}|\bm{q},\bm{\epsilon},\bm{x}_{0})\right\rangle_{\bm{q},\bm{\epsilon},\bm{x}_{0}}=\lim_{N\rightarrow\infty}\lim_{\beta\rightarrow\infty}\frac{1}{\beta N}\Big{\langle}\log Z_{N}(\beta|\bm{q},\bm{\epsilon},\bm{x}_{0})\Big{\rangle}_{\bm{q},\bm{\epsilon},\bm{x}_{0}}\ .$ (13)
From a practical standpoint, the above expression looks rather inconvenient, as it requires averaging over the logarithm of the partition function $Z_{N}$. Such a rather cumbersome task is usually circumvented via the so-called replica trick, which effectively lets us compute the average of the logarithm as the logarithm of an average. This is achieved by first converting the average over $\log Z_{N}$ into an average over $r$ replicas of the system—that is, the factorization of $r$ identical partition functions—through the formal identity
$\langle\log Z_{N}\rangle=\lim_{r\rightarrow 0}\frac{\langle Z_{N}^{r}\rangle-1}{r}\ .$
When using replicas, $r$ is assumed to be an integer number throughout all calculations, and restored to be a real number only at the end by taking the above limit.
All in all, the optimization problem in Eq. (12) can be solved as
$\lim_{N\rightarrow\infty}\frac{1}{N}\left\langle\max_{\bm{s}\geq 0}\ U_{N}(\bm{s}|\bm{q},\bm{\epsilon},\bm{x}_{0})\right\rangle_{\bm{q},\bm{\epsilon},\bm{x}_{0}}=\lim_{N\rightarrow\infty}\lim_{\beta\rightarrow\infty}\lim_{r\rightarrow 0}\frac{1}{\beta Nr}\log\Big{\langle}Z_{N}^{r}(\beta|\bm{q},\bm{\epsilon},\bm{x}_{0})\Big{\rangle}_{\bm{q},\bm{\epsilon},\bm{x}_{0}}=h(\bm{\theta}^{\star})\ ,$ (14)
where the last equality comes from the fact that the replica trick allows to identify a small set of order parameters $\bm{\theta}=(\theta_{1},\ldots,\theta_{k})$ such that
$\Big{\langle}Z_{N}^{r}(\beta|\bm{q},\bm{\epsilon},\bm{x}_{0})\Big{\rangle}_{\bm{q},\bm{\epsilon},\bm{x}_{0}}\approx\int\mathrm{d}\bm{\theta}\ \mathrm{e}^{\beta Nrh(\bm{\theta})}$ (15)
In Appendix B we provide a detailed derivation of the solution to the above optimization problem, including the identification of the corresponding order parameters. Let us emphasize here that identifying a system’s order parameters in a replica trick calculation is more of an art than a science, as there is no straightforward ‘algorithmic’ procedure to do that. However, identifying a system’s order parameters simplifies its description quite considerably. As a matter of fact, one should not lose sight of the fact that, when taking the thermodynamic limit, one is dealing with infinitely many variables (infinitely many firms and goods, in our case). Yet, when the order parameters have been identified, one ends up—through Eq. (14)—with a complete description of the system based on just $k$ variables, where $k$ is usually a small number (in our case, we will see that $k=6$).
Appendix B The optimization problem
@@
$\left.\times\exp\left[\beta\sum_{a=1}^{r}U_{N}\left(\bm{x}_{a}\right)+\mathrm{i}\sum_{a=1}^{r}\sum_{c=1}^{C}\hat{x}_{a}^{c}\left(x_{a}^{c}-x_{0}^{c}\right)\right]\right$.
$\left.\times\exp\left[-\frac{1}{2C}\sum_{n=1}^{N}\sum_{c=1}^{C}\left(\hat{z}_{n}-\sum_{a=1}^{r}s_{n,a}\hat{x}_{a}^{c}\right)^{2}+\mathrm{i}\sqrt{\lambda}w\sum_{n=1}^{N}\hat{z_{n}}+\mathrm{i}\epsilon_{0}\sum_{n=1}^{N}\hat{z}_{n}\right]\right\rangle_{w}\ .$
In the above expression, the average $\langle\cdots\rangle_{w}$ comes from the Hubbard-Stratonovich transformation
$\exp\left(-\frac{ax^{2}}{2}\right)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathrm{d}w\exp\left(-\frac{w^{2}}{2}+\mathrm{i}\sqrt{a}xw\right)=\Big{\langle}\exp\left(\mathrm{i}\sqrt{a}xw\right)\Big{\rangle}_{w}\ ,$
which linearizes the term proportional to $\hat{z}^{2}$ at the cost of introducing an auxiliary Gaussian standard variable $w\sim\mathcal{N}(0,1)$. The prefactor $\left[\sqrt{2\pi\left(\lambda+1\right)}\exp\left(\epsilon_{0}^{2}/2\left(\lambda+1\right)\right)\right]^{N}$ comes from integrating out the denominator in Eq. (8) with respect to $\epsilon$ (i.e., with respect to the pdf in Eq. (7)).
@@
$1$ $=$ $N\int_{-\infty}^{\infty}\mathrm{d}\omega_{ab}\ \delta\left(N\omega_{ab}-\sum_{n=1}^{N}s_{n,a}s_{n,b}\right)=N\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega_{ab}\mathrm{d}\hat{\omega}_{ab}}{2\pi\mathrm{i}}\exp\left[\hat{\omega}_{ab}\left(N\omega_{ab}-\sum_{n=1}^{N}s_{n,a}s_{n,b}\right)\right]$
$1$ $=$ $N\int_{-\infty}^{\infty}\mathrm{d}k_{a}\ \delta\left(Nk_{a}-\sum_{n=1}^{N}\hat{z}_{n}s_{n,a}\right)=N\int_{-\infty}^{\infty}\frac{\mathrm{d}k_{a}\mathrm{d}\hat{k}_{a}}{2\pi\mathrm{i}}\exp\left[\hat{k}_{a}\left(Nk_{a}-\sum_{n=1}^{N}\hat{z}_{n}s_{n,a}\right)\right]\ ,$
@@
[…]
@@
[…]
@@
Posing $p=-\hat{\kappa}$ and $\sigma=\sqrt{\hat{\gamma}-\left(1+\lambda\right)\hat{\kappa}^{2}}$, Eq. (B) can be rewritten as the following expression:
[…]
$+$ $\frac{1}{\nu}\left\langle\max_{s\geq 0}\left[u\left(x\right)-\frac{1}{2\chi}\left(x-x_{0}+t\sqrt{\nu\Omega}+\frac{\kappa}{1+\lambda}\right)^{2}\right]\right\rangle_{t,x_{0}}\ ,$
and the saddle point condition (see Eq. (21)) provides us with a set of equations for the order parameters that fully specify the solution of our optimization problem:
$\Omega$ $=$ $\left\langle\left(s^{\star}\right)^{2}\right\rangle_{t}$ (21)
@@
Appendix C Statistical properties at equilibrium
In this section we derive the statistical properties of the optimal scales of production $s^{\star}$ and consumption $x^{\star}$ at equilibrium. Starting from the former, we see from the third term in Eq. (B) that it must be the solution of the following equation